OpenStudy (anonymous):

insert =, <, or > to make a true statement: -1/18 ...... -2/11

OpenStudy (anonymous):

-1/18 vs -2/11 -1/18 = -11/198 -2/11 = -36/198 Therefore, -1/18 > -2/11

OpenStudy (anonymous):

thankyou((:

OpenStudy (amistre64):

cut a pie into 18 pieces; 1 pieces is small cut another pie into 11 pieces; you see that theya re bogger piees and you get 2 of them now; 2/11 is bigger then 1/18; but when we negate these; the bigger one becomes smaller -1/8 > -2/11

OpenStudy (anonymous):

okay. thank you!!(((: